Mr. Smith has $100.00 to buy 100 animals with. Cows cost $10.00 each, pigs cost $3.00 each and chickent cost $.50 each. How many can he buy with exactly $100.00.
I have this equation so far, but where do I go from here?
10x + 3y + .5z = 100

Found 2 solutions by stanbon, AnlytcPhil:
Answer by stanbon(75887)

I think there is more information involved in this
problem. As you stated it he could buy 10 cows or
200 chickens or many other combinations of cows,
pigs, and chickens. Please check the problem again.
Cheers,
Stan H.

Answer by AnlytcPhil(1806)

Mr. Smith has $100.00 to buy 100 
animals with. Cows cost $10.00 each, 
pigs cost $3.00 each and chickens 
cost $.50 each. How many can he buy 
with exactly $100.00.  I have this 
equation so far, but where do I go 
from here? 

10x + 3y + .5z = 100

Multiply through by 10 to clear decimal:

100x + 30y + 5z = 1000

Divide by 5

20x + 6y + z = 200
   x + y + z = 100

19x + 5y = 100

19x = 100 - 5y

19x = 5(20 - y)

  x = 5(20 - y)/19

Since x is a positive integer,
the 19 must divide evenly into
5(20 - y).  the 19 will not
divide evenly into the 5, so
19 must divide into 20 - y.

20 - y is at most 19, since
y must be at least 1.  The
only integer at least 19
which is divisible by 19 is
19 itself.  Therefore

20 - y = 19
    -y = -1
     y = 1 
 
x = 5(20 - y)/19
x = 5(20 - 1)/19
x = 5(19)/19
x = 5

x + y + z = 100
5 + 1 + z = 100
    6 + z = 100
        z = 94

Therefore he bought
5 cows, 1 pig, 
and 94 chickens.

Edwin